while finding the amount of seed needed to plant his three square wheat fields, Hank observed that the side of one field was 1 kilometer longer than the side of the smallest field and the side of the longest was 3 kilometers longer than the side of the smallest field. if the area of the three fields is 38 square kilometers, then what is the area of each field?
Answer by checkley77(12844)

X^2+(X+1)^2+(X+3)^2=38
X^2+X^2+2X+1+X^2+6X+9=38
3X^2+8X+10-38=0
3X^2+8X-28=0
(3X+14)(X-2)=0
X-2=0
X=2 ANS.
PROOF:
(2)^2+(2+1)^2+(2+3)^2=38
4+3^2+5^2=38
4+9+25=38
38=38